OpenStudy (anonymous):

simplify \[2 \sqrt{108}\]

OpenStudy (anonymous):

108 = 12*9 So, \[2\sqrt{108}=2\sqrt{3*3*2*2*3}=2*3*2\sqrt{3}=12\sqrt{3}\]